OpenStudy (anonymous):

derivative of e^(cos^-1(x))

OpenStudy (anonymous):

d/dx(e^(cos^(-1)(x))) = -e^(cos^(-1)(x))/sqrt(1-x^2)

OpenStudy (anonymous):

y=e^(arccos(x)) take antilog on both side lny=arccos(x) 1/y y'=1/sqrt(1-x^2) so y'=e^(arccos(x)) /sqrt(1-x^2)